Standard form of a quadratic function (also called (h,k) form)
*change the function to (h,k) form: y=2x^2 +4x +4

Answer by stanbon(57967)

change the function to (h,k) form: y=2x^2 +4x +4
-------------
Form y = a(x-h)^2 + k
---------------------------------
y=2x^2 +4x +4
Complete the square on the x-terms.
y = 2(x^2+2x+1) + 4-2
-----
y = 2(x+2)^2 + 2
===================
Cheers,
Stan H.
===================